# DEFAME: Dynamic Evidence-based FAct-checking with Multimodal Experts

### v2

@@ 4.7 Failure Analysis
We analyzed 20 VERITE cases where DEFAME mislabeled a sample and identified three common failure modes. (1) Premature Judgment, the most frequent one, occurs when DEFAME commits to a verdict before gathering sufficient evidence. Often, it predicts [Uncaptioned image]Supported as soon as it finds any image matching the claim text, even if it is not the one in the claim. (2) Failed Evidence Retrieval accounts for 25% of errors, where DEFAME exhausts all three retrieval attempts but fails to find enough evidence. This happens mainly when evidence is inaccessible (e.g., when API access is required) or embedded in a video. (3) Wrong Ground Truth: In three cases, our inspection revealed annotation errors, meaning DEFAME’s prediction was actually correct. In one instance, DEFAME “over-summarized” the evidence, omitting critical details.
@@ 5 Discussion
Credibility of External Evidence: Our reliance on search engines introduces the risk of incorporating unreliable information, as search results may include leaked or biased content (Chrysidis et al., 2024). While some approaches assess source credibility using third-party ratings  (Zhou et al., 2024), our framework relies on diversification and the search engine’s internal trustworthiness checks121212blog.google/products/search/how-google-delivers-reliable-information-search/. Introducing external ratings could strengthen verification rigor.
@@ 6 Conclusion
We presented DEFAME, a comprehensive zero-shot framework for multimodal fact-checking that integrates MLLMs with external tools to address the limitations of traditional automated fact-checking approaches. Our framework explainably grounds its analysis in verifiable data by combining MLLM-driven reasoning with external multimodal evidence sources, geolocation, reverse image search, and dynamic web searches. We demonstrated that DEFAME achieves new state-of-the-art results across multiple diverse benchmarks, including an own leakage-free benchmark, highlighting its capability to navigate unseen real-world claims. While limitations remain, our system represents a significant advancement in the field, providing a versatile foundation for future developments in automated multimodal fact-checking.
7 Impact Statement
In order to scalably debunk the masses of mis- and disinformation—which indisputably pose a strong threat to social cohesion—automating fact-checking is inevitable. Although the results do not sufficiently warrant real-world application yet, DEFAME’s introduction marks a pivotal step in addressing that challenge.
DEFAME contributes shifting the paradigm toward explainable and verifiable AI systems. Its ability to provide structured, evidence-backed justifications ensures that AFC does not become a black-box authority but rather an augmentation tool that strengthens public resilience against misinformation. By embedding transparency and iterative verification into its core process, DEFAME helps counter the skepticism toward AI-driven moderation while simultaneously reducing the burden on human fact-checkers.
However, the widespread deployment of such systems also raises ethical and societal concerns. Recently, in public discussion, professional fact-checkers have widely faced accusations of being biased and, therefore, unreliable. Automated fact-checking, even when well-calibrated, can also shape narratives by prioritizing certain sources or by amplifying institutional biases inherent in retrieval mechanisms or parametric knowledge. A recent example for governmental censorship in LLMs is the DeepSeek LLM family. Current research efforts strive to de-bias LLMs, which DEFAME will automatically benefit from thanks to its backbone-agnostic design. Within the strongly polarized public discussions, AFC systems may emerge as a bipartisan information source. Their reliability inherently depends on the quality of the backbone, which is difficult to rigorously analyze.
@@ 7 Impact Statement
8 Acknowledgments
@@
The appendix provides an in-depth extension of the main paper, showcasing additional examples, analyses, and details that complement the findings discussed. Section A outlines excluded domains to ensure a fair and realistic evaluation setting. We include information on compute and cost in Section B. Details about our adaptations for the AVeriTeC challenge are given in Section C. A short analysis of the confusions committed by the CoT variant and DEFAME is shown in Section D. Section E includes the prompts used in DEFAME. Section F identifies wrongly annotated samples in the VERITE dataset. Section G discusses cases with outdated ground truth in the MOCHEG dataset and their implications. Section H specifies our human evaluation of the fact-checking report. In Section  I, we analyze the justifications generated by DEFAME and scrutinize current evaluation metrics. We present our ClaimReview2024+ dataset in Section J. Section K illustrates an example fact-checking report produced by DEFAME. Lastly, Section L delves into failure cases, highlighting common pitfalls and their underlying causes.
Appendix A Excluded Domains
@@
Model Variant
Time min / Claim
LLM API Cost ¢ / Claim
Search API Cost ¢ / Claim
# Search Queries Searches / Claim
# Input Tokens Processed Tokens / Claim
# Output Tokens Generated Tokens / Claim
DEFAME:
- GPT-4o 4:14 13 $<$1 2.2 48K 921
- GPT-4o mini 3:13 $<$1 $<$1 3.1 54K 1300
- LLaVA-1V 1:23 - $<$1 0.9 14K 794
w/o Geolocation 3:12 11 $<$1 2.3 40K 910
w/o Reverse Search 2:37 8 $<$1 2.3 28K 828
w/o Image Search 2:57 11 $<$1 2.2 40K 909
w/o Web Search 2:30 10 $<$1 2.2 37K 761
Single Turn 2:47 10 $<$1 2.0 38K 858
Static Actions 5:11 14 $<$1 5.2 70K 1100
w/o Develop 1:56 10 $<$1 2.1 42K 710
Unimodal Develop 3:15 10 $<$1 2.3 41K 865
Pure GPT-4o 0:06 $<$1 0 0.0 2.2K 67
Chain-of-Thought 0:09 $<$1 0 0.0 2.3K 177
Table 7: Usage and computational cost statistics of DEFAME and all considered variants on VERITE. Time values are rough estimates and object to high variance due to hardware usage by other processes.
Appendix B Details on Compute and Cost
As expected, the naive baselines (Pure GPT-4o and CoT) require the least amount of resources. But, as discussed in Section 4, this is due to the absence of any external evidence and comes at the cost of lower accuracy. DEFAME with GPT-4o mini and LLaVA-1V as the backbone MLLM runs faster and cheaper than DEFAME with GPT-4o, but yields decreased accuracy as well. Surprisingly, LLaVA-1V fact-checks faster than the GPT models. We found the reason to lie in LLaVA-1V’s inability to correctly format the proposed actions, yielding a smaller number of executed actions and, thus, shorter fact-checks overall.
Appendix C Adaptation to the AVeriTeC Challenge
@@
To align with this evaluation, we introduce an extension called DEFAME-QA. This adaptation begins by generating 10 key questions designed to probe the claim’s veracity. The Planner then proposes targeted search queries for each question and applies the Web-Search tool to retrieve up to 5 relevant search results. Using the retrieved evidence, the LLM backbone attempts to answer the questions systematically. Finally, the system synthesizes the answers into a coherent verdict, ensuring the reasoning is grounded in the collected evidence. The resulting QA pairs are evaluated using the AVeriTeC Score, showcasing DEFAME-QA’s alignment with the benchmark’s evaluation criteria while maintaining its structured, evidence-driven methodology.
Appendix D Model Confusions
@@
A closer look at Figures 7(a), 7(b) and 7(c) reveals that GPT-4o with and without Chain-of-Thought overpredicts [Uncaptioned image]OOC while DEFAME’s confusions are more balanced. Surprisingly, incorporating Chain-of-Thought prompting hurts the performance on MOCHEG (see Figures 7(d), 7(e) and Table  3). According to the confusion matrices the introduction of Chain-of-Thought makes the model more unsure, leaning towards [Uncaptioned image]NEI compared to the Pure GPT-4o. In contrast, DEFAME predicts too confidently even when the ground truth implies insufficient information. A qualitative analysis of the failure cases reveals that, in several cases, the ground truth explanation is no longer up-to-date (see Sections 4.7 and Appendix F).
On AVeriTeC (Figures 7(g), 7(h), and 7(i)), we observe the opposite behavior with the GPT-4o variants overpredicting [Uncaptioned image]NEI compared to our framework. Lastly, DEFAME’s false predictions on the ClaimReview2024+ dataset appear balanced in Figure 4 while the baselines lack the necessary evidence to make correct veracity predictions, very often defaulting to [Uncaptioned image]NEI.
(a) VERITE - Pure GPT-4o
(b) VERITE - Chain-of-Thought
(c) VERITE - DEFAME
(d) MOCHEG - Pure GPT-4o
(e) MOCHEG - Chain-of-Thought
(f) MOCHEG - DEFAME
(g) AVeriTeC - Pure GPT-4o
(h) AVeriTeC - Chain-of-Thought
(i) AVeriTeC - DEFAME
Figure 7: Confusion matrices for Pure GPT-4o, Chain-of-Thought, and DEFAME across the VERITE, MOCHEG, and AVeriTeC datasets.
Appendix E LLM Prompts
Each stage of the DEFAME framework is guided by a tailored prompt. These prompts are constructed from prompt templates, where values enclosed in [] serve as placeholders. During execution, these placeholders are dynamically replaced with the corresponding variables. This process ensures that the prompt is specific to the current task and context, as illustrated in E.1 and E.2.
@@ Appendix E LLM Prompts
E.1 Plan Prompt Template
Instructions The available knowledge is insufficient to assess the Claim. Therefore, propose a set of actions to retrieve new and helpful evidence. Adhere to the following rules: The actions available are listed under Valid Actions, including a short description for each action. No other actions are possible at this moment. For each action, use the formatting as specified in Valid Actions. Include all actions in a single Markdown code block at the end of your answer. Propose as few actions as possible but as many as needed. Do not propose similar or previously used actions.
[Extra Rules]
[Valid Actions]
[Examples]
[Record]
Your Actions:
@@ E.1 Plan Prompt Template
E.2 Finalized Plan Prompt
Instructions The available knowledge is insufficient to assess the Claim. Therefore, propose a set of actions to retrieve new and helpful evidence. Adhere to the following rules: The actions available are listed under Valid Actions, including a short description for each action. No other actions are possible at this moment. For each action, use the formatting as specified in Valid Actions. Include all actions in a single Markdown code block at the end of your answer. Propose as few actions as possible but as much as needed. Do not propose similar or previously used actions. Consider Both Modalities Equally: Avoid focusing too much on one modality at the expense of the other, but always check whether the text claim is true or false. Compare Image and Caption: Verify the context of the image and caption.
Valid Actions: geolocate: Determine the country where an image was taken by providing an image ID. reverse_search: Perform a reverse image search on the web for similar images. web_search: Run an open web search for related webpages. image_search: Retrieve related images for a given query.
Examples: geolocate(<image:k>) reverse_search(<image:k>) web_search("New Zealand Food Bill 2020") image_search("China officials white suits carry people")
Record: Claim: “<image:1232>Image of a bus powered by compressed natural gas, bursting into flames in Italy.”
Your Actions:
E.3 Summarize Prompt Template
Instructions In order to find evidence that helps your fact-check, you just ran a web search, which yielded a Search Result. Your task right now is to summarize the Search Result. What to include: Information that might be useful for the fact-check (see Record). Relevant images (refer to images by inserting their reference in the format <image:k>). If available: the release date as well as the author or the publisher (e.g., the media company) of the search result. Do NOT include: Advertisements. Any other information unrelated to the Record or the Claim. Additional Rules: Do not add any additional information besides the information in the Search Result. If the Search Result doesn’t contain any relevant information for the fact-checking work, simply print one word in capital letters: NONE. Keep your writing style consistent with the provided Examples. Try to filter out relevant information even if the Search Result is in a different language.
[Examples]
[Record]
[Search_Result]
Your Summary:

### v3

@@ 5.7 Failure Analysis
We analyzed 119 VERITE and ClaimReview2024+ instances that were mislabeled by DEFAME and identified five common failure modes. (1) Label Ambiguity, the most frequent one, occurs when DEFAME mixes up similar labels like [Uncaptioned image]Refuted and [Uncaptioned image]Misleading. (2) Missed Evidence accounts for errors where DEFAME exhausts all three retrieval attempts but fails to find enough evidence. This happens often when evidence is inaccessible (e.g., when proprietary API access is required) or embedded in a video. (3) Reasoning Errors are mistakes that happen during natural language inference, e.g., mixing up numbers. (4) Premature Judgment occurs when DEFAME commits to a verdict before gathering sufficient evidence. For example, it predicts [Uncaptioned image]Supported as soon as it finds any image matching the claim text, even if it is not the one in the claim. (5) Wrong Ground Truth: In several cases, our inspection revealed annotation errors, meaning DEFAME’s prediction was actually correct. Appendix I shows instances with wrong ground truth labels.
@@ 6 Discussion
Credibility of External Evidence: Our reliance on search engines introduces the risk of incorporating unreliable information, as search results may include leaked or biased content (Chrysidis et al., 2024). While some approaches assess source credibility using third-party ratings  (Zhou et al., 2024), our framework relies on diversification and the search engine’s internal trustworthiness checks [^9] https://blog.google/products/search/how-google-delivers-reliable-information-search/. Introducing external ratings could strengthen verification rigor.
@@ 7 Conclusion
We presented DEFAME, a comprehensive zero-shot framework for multimodal fact-checking that integrates MLLMs with external tools to address the limitations of traditional automated fact-checking approaches. Our framework explainably grounds its analysis in verifiable data by combining MLLM-driven reasoning with external multimodal evidence sources, geolocation, reverse image search, and dynamic web searches. We demonstrated that DEFAME achieves new state-of-the-art results across multiple diverse benchmarks, including an own leakage-free benchmark, highlighting its generality and capability to navigate unseen real-world claims. While limitations remain, our system represents a significant advancement in the field, providing a versatile foundation for future developments in automated multimodal fact-checking.
Impact Statement
To scalably debunk the masses of mis- and disinformation—which indisputably pose a substantial threat to social cohesion—automating fact-checking is inevitable. Although the results do not sufficiently warrant real-world application yet, DEFAME’s introduction is a step in addressing that challenge.
DEFAME contributes to shifting the paradigm toward explainable and verifiable AI systems. Its ability to provide structured, evidence-backed justifications ensures that AFC does not become a black-box authority but rather an augmentation tool that strengthens public resilience against misinformation. By embedding transparency and iterative verification into its core process, DEFAME helps counter the skepticism toward AI-driven moderation while simultaneously reducing the burden on human fact-checkers.
However, the widespread deployment of such systems also raises ethical and societal concerns. Recently, in public discussions, professional fact-checkers have widely faced accusations of being biased (CNN, 2025) and, therefore, unreliable. Automated fact-checking, even when well-calibrated, can also shape narratives by prioritizing certain sources or by amplifying institutional biases inherent in retrieval mechanisms or parametric knowledge. Current research efforts strive to de-bias LLMs (Gallegos et al., 2024), which DEFAME will automatically benefit from thanks to its backbone-agnostic design. Within the strongly polarized public discussions, AFC systems may emerge as a nonpartisan information source. Their reliability inherently depends on the quality of the backbone, which is difficult to rigorously analyze.
@@
Acknowledgments
We thank Marcus Kornmann for his diligent assistance in conducting this research.
@@ Acknowledgments
For compute, we gratefully acknowledge support from the hessian.AI Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant no. 01IS22091) and the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation, grant no. S-DIW04/0013/003).
@@
The appendix provides an in-depth extension of the main paper, showcasing additional examples, analyses, and details that complement the findings discussed. The data leakage problem of existing AFC benchmarks is discussed in more detail in Section A. In Section B, the definitions of the four ClaimReview2024+ classes are presented. We include information on compute and cost in Section C. Section D outlines excluded domains to ensure a fair and realistic evaluation setting. Details about our adaptations for the AVeriTeC challenge are given in Section E. To enhance clarity, we formalize the iterative structure and processing stages of DEFAME in Appendix F. A short analysis of the confusions committed by the CoT variant and DEFAME is shown in Section G. Section H includes the prompts used in DEFAME. Section I identifies wrongly annotated samples in the VERITE dataset. Section J provides a high-level derivation showing the equivalence of micro-F1 and accuracy in single-label multiclass settings, clarifying a common source of confusion in prior evaluations. Section K discusses cases with outdated ground truth in the MOCHEG dataset and their implications. Section L specifies our human evaluation of the fact-checking report. In Section  M, we analyze the justifications generated by DEFAME and scrutinize current evaluation metrics. Section N illustrates example fact-checking reports produced by DEFAME. Lastly, Section O delves into failure cases, highlighting common pitfalls and their underlying causes.
Appendix A The Data Leakage Issue
Most AFC benchmarks, particularly AVeriTeC, VERITE, and MOCHEG, are (largely) composed of real-world claims sourced from actual fact-checking articles on the web. The covered time spans vary:
  • AVeriTeC claims stem from the time until December 2021.
  • VERITE covers claims between January 2001 and January 2023.
  • MOCHEG contains claims up until May 2022.
Importantly, all time ranges lie before GPT-4o’s knowledge cutoff date, which is October 2023. Since the claims are publicly fact-checked, it is highly likely that GPT-4o was pretrained on large portions of the claims and verdicts contained in these benchmarks, a matter of data leakage. Therefore, we introduced ClaimReview2024+, which consists of claims, the fact-checks of which were published only after the knowledge cutoff date.
Appendix B ClaimReview2024+ Class Definitions
ClaimReview2024+ uses four classes, defined as follows:
  • [Uncaptioned image]Refuted: “A claim is considered refuted when the evidence contradicts the claim.”
  • [Uncaptioned image]Supported: “The claim is accurate based on evidence.”
  • [Uncaptioned image]Misleading: “The claim is misleading or requires additional context.”
  • [Uncaptioned image]NEI: “The claim does not have enough information to be verified.”
For all datasets, the class definitions were provided to DEFAME at the judgment stage.
Appendix C Details on Compute and Cost
Method
Time min / Claim
LLM API Cost ¢ / Claim
Search API Cost ¢ / Claim
# Search Queries Searches / Claim
# Input Tokens Processed Tokens / Claim
# Output Tokens Generated Tokens / Claim
DEFAME GPT-4o 4:14 13 $<$1 2.2 48K 921
GPT-4o mini 3:13 $<$1 $<$1 3.1 54K 1300
LLaVA-1V 1:23 - $<$1 0.9 14K 794
Llama 4 18:02 - $<$1 2.8 21K 1325
DEFAME Ablations w/o Geolocation 3:12 11 $<$1 2.3 40K 910
w/o Reverse Search 2:37 8 $<$1 2.3 28K 828
w/o Image Search 2:57 11 $<$1 2.2 40K 909
w/o Web Search 2:30 10 $<$1 2.2 37K 761
Single Turn 2:47 10 $<$1 2.0 38K 858
w/o Planning 5:11 14 $<$1 5.2 70K 1100
w/o Develop 1:56 10 $<$1 2.1 42K 710
Unimodal Develop 3:15 10 $<$1 2.3 41K 865
Baselines GPT-4o 0:06 $<$1 0 0.0 2.2K 67
GPT-4o CoT 0:09 $<$1 0 0.0 2.3K 177
Table 5: Usage and computational cost statistics of DEFAME and all considered variants (ablations and baselines) on VERITE. Time values are rough estimates and are subject to high variance due to hardware usage by other processes.
As expected, the naive baselines (Pure GPT-4o and CoT) require the least amount of resources. But, as discussed in Section 5, this is due to the absence of any external evidence and comes at the cost of lower accuracy. DEFAME with GPT-4o mini and LLaVA-1V as the backbone MLLM runs faster and cheaper than DEFAME with GPT-4o, but yields decreased accuracy as well. Surprisingly, LLaVA-1V fact-checks faster than the GPT models. We found the reason to lie in LLaVA-1V’s inability to correctly format the proposed actions, yielding a smaller number of executed actions and, thus, shorter fact-checks overall. Llama 4 Scout masters action formatting and is cheaper than the GPT-4o models but—surprisingly—takes drastically more time. We have not taken any acceleration measures beyond basic LLM installation, which may leave room for time improvement.
Appendix D Excluded Domains
@@
Appendix E Adaptation to the AVeriTeC Challenge
@@
To align with this evaluation, we introduce an extension called InFact. This adaptation begins by generating $10$ key questions designed to probe the claim’s veracity. The Planner then proposes targeted search queries for each question and applies the Web Search tool to retrieve up to $5$ relevant search results. Using the retrieved evidence, the LLM backbone attempts to answer the questions systematically. Finally, the system synthesizes the answers into a coherent verdict, ensuring the reasoning is grounded in the collected evidence. The resulting QA pairs are evaluated using the AVeriTeC Score, showcasing InFact ’s alignment with the benchmark’s evaluation criteria while maintaining its structured, evidence-driven methodology.
Appendix F Formal Representation of DEFAME
We include a formalization of the DEFAME pipeline to clarify the role of each stage and its iterative structure.
Let $\mathcal{T}$ and $\mathcal{I}$ denote the spaces of text and images, respectively. Define $\mathcal{M}:=(\mathcal{T}\cup\mathcal{I})^{*}$ as the space of multimodal sequences, and let $\mathcal{Y}$ denote the space of verdict labels. Then, DEFAME is a function
:×$\mathcal{F}:\mathcal{M}\to\mathcal{M}\times\mathcal{Y},\quad\mathcal{F}(c)=(R_{\text{out}},y_{\text{out}}),$
where, given a claim $c\in\mathcal{M}$, the output consists of a report $R_{\text{out}}$ containing the full fact-check and a predicted verdict $y_{\text{out}}$. DEFAME proceeds iteratively up to $N$ steps. We can denote each iteration with $\mathcal{F}_{\text{iter}}:\mathcal{M}\rightarrow\mathcal{M}\times\mathcal{Y}$, so that
$(R^{(i+1)},y^{(i+1)}):=\mathcal{F}_{\text{iter}}(R^{(i)}),$
i.e., an (incomplete) report $R^{(i)}\in\mathcal{M}$ gets extended with new actions, evidence, and elaboration, resulting in report $R^{(i+1)}$ and intermediate verdict $y^{(i+1)}$. We can decompose $\mathcal{F}_{\text{iter}}$ into the five individual pipeline stages
$\mathcal{F}_{\text{iter}}:=\mathcal{S}_{5}\circ\mathcal{S}_{4}\circ\mathcal{S}_{3}\circ\mathcal{S}_{2}\circ\mathcal{S}_{1},$
where each stage can be described as follows:
  1. 1.
    Planning ($\mathcal{S}_{1}$): Select actions $A\subseteq\mathcal{A}$ based on the current report $R^{(i)}$.
  2. 2.
    Execution ($\mathcal{S}_{2}$): Retrieve evidence $E:=\{\tau(a)\mid a\in A\}$, where $\tau$ executes the tool action $a$.
  3. 3.
    Summarization ($\mathcal{S}_{3}$): $R_{1}^{(i)}:=\sigma(E,R^{(i)})$, where $\sigma$ summarizes $E$ in context and appends it to the report.
  4. 4.
    Develop ($\mathcal{S}_{4}$): $R_{2}^{(i)}:=\mathcal{S}_{4}(R_{1}^{(i)})$, where $\mathcal{S}_{4}$ generates structured reasoning and expands the report.
  5. 5.
    Verdict Prediction ($\mathcal{S}_{5}$): $(R_{3}^{(i)},y^{(i)}):=\mathcal{S}_{5}(R_{2}^{(i)})$, returning an updated report $R_{3}^{(i)}\in\mathcal{M}$ and a verdict $y^{(i)}\in\mathcal{Y}$.
Let $i^{*}:=\min\{i\leq N\mid y^{(i)}\neq\text{NEI}\text{ or }i=N\}$, then the final outputs are $y_{\text{out}}:=y^{(i^{*})}$ and $R_{\text{out}}:=\mathcal{S}_{6}(R^{(i^{*})})$, where $\mathcal{S}_{6}$ denotes the justification stage that appends a rationale to the final report.
This formal view captures the iterative and modular nature of DEFAME, highlighting how evidence is retrieved, processed, and transformed into a final verdict and explanation.
Appendix G DEFAME Confusions
@@
A closer look at Figures 8(g), 8(h) and 8(i) reveals that GPT-4o with and without Chain-of-Thought overpredicts [Uncaptioned image]OOC while DEFAME’s confusions are more balanced. Surprisingly, incorporating Chain-of-Thought prompting hurts the performance on MOCHEG (see Figures 8(d), 8(e) and Table  3). According to the confusion matrices the introduction of Chain-of-Thought makes the model more unsure, leaning towards [Uncaptioned image]NEI compared to the Pure GPT-4o. In contrast, DEFAME predicts too confidently even when the ground truth implies insufficient information. A qualitative analysis of the failure cases reveals that, in several cases, the ground truth explanation is no longer up-to-date (see Sections 5.7 and Appendix I).
On AVeriTeC (Figures 8(a), 8(b), and 8(c)), we observe the opposite behavior with the GPT-4o variants overpredicting [Uncaptioned image]NEI compared to our framework. Lastly, DEFAME’s false predictions on the ClaimReview2024+ dataset appear balanced in Figure 4 while the baselines lack the necessary evidence to make correct veracity predictions, very often defaulting to [Uncaptioned image]NEI.
(a) AVeriTeCGPT-4o
(b) AVeriTeCGPT-4o CoT
(c) AVeriTeCDEFAME
(d) MOCHEG – GPT-4o
(e) MOCHEG – GPT-4o CoT
(f) MOCHEG – DEFAME
(g) VERITE – GPT-4o
(h) VERITE – GPT-4o CoT
(i) VERITE – DEFAME
Figure 8: Confusion matrices for GPT-4o, GPT-4o CoT, and DEFAME across the AVeriTeC, MOCHEG, and VERITE datasets.
Appendix H Prompts Used in DEFAME
Each stage of the DEFAME framework is guided by a tailored prompt. These prompts are constructed from prompt templates, where values enclosed in [] serve as placeholders. During execution, these placeholders are dynamically replaced with the corresponding variables. This process ensures that the prompt is specific to the current task and context, as illustrated in H.1 and H.2.
@@ Appendix H Prompts Used in DEFAME
H.1 Plan Prompt Template
Instructions The available knowledge is insufficient to assess the Claim. Therefore, propose a set of actions to retrieve new and helpful evidence. Adhere to the following rules: The actions available are listed under Valid Actions, including a short description for each action. No other actions are possible at this moment. For each action, use the formatting as specified in Valid Actions. Include all actions in a single Markdown code block at the end of your answer. Propose as few actions as possible but as many as needed. Do not propose similar or previously used actions. [Extra Rules] [Valid Actions] [Examples] [Record] Your Actions:
@@ H.1 Plan Prompt Template
H.2 Finalized Plan Prompt
Instructions The available knowledge is insufficient to assess the Claim. Therefore, propose a set of actions to retrieve new and helpful evidence. Adhere to the following rules: The actions available are listed under Valid Actions, including a short description for each action. No other actions are possible at this moment. For each action, use the formatting as specified in Valid Actions. Include all actions in a single Markdown code block at the end of your answer. Propose as few actions as possible but as much as needed. Do not propose similar or previously used actions. Consider Both Modalities Equally: Avoid focusing too much on one modality at the expense of the other, but always check whether the text claim is true or false. Compare Image and Caption: Verify the context of the image and caption. Valid Actions: geolocate: Determine the country where an image was taken by providing an image ID. reverse_search: Perform a reverse image search on the web for similar images. web_search: Run an open web search for related webpages. image_search: Retrieve related images for a given query. Examples: geolocate(<image:k>) reverse_search(<image:k>) web_search("New Zealand Food Bill 2020") image_search("China officials white suits carry people") Record: Claim: “<image:1232> Image of a bus powered by compressed natural gas, bursting into flames in Italy.” Your Actions:
H.3 Summarize Prompt Template
Instructions In order to find evidence that helps your fact-check, you just ran a web search, which yielded a Search Result. Your task right now is to summarize the Search Result. What to include: Information that might be useful for the fact-check (see Record). Relevant images (refer to images by inserting their reference in the format <image:k>). If available: the release date as well as the author or the publisher (e.g., the media company) of the search result. Do NOT include: Advertisements. Any other information unrelated to the Record or the Claim. Additional Rules: Do not add any additional information besides the information in the Search Result. If the Search Result doesn’t contain any relevant information for the fact-checking work, simply print one word in capital letters: NONE. Keep your writing style consistent with the provided Examples. Try to filter out relevant information even if the Search Result is in a different language. [Examples] [Record] [Search_Result] Your Summary: